OpenStudy (anonymous):

if 4x+8y-8=10 what is the value of x+2y

OpenStudy (anonymous):

9/2

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

4x+8y=18 dvide by 4 on both side x+2y=9/2

OpenStudy (anonymous):

9/2

jhonyy9 (jhonyy9):

4x+8y=18 4(x+2y)=18 x+2y=18/4 x+2y=9/2